The Target Copy Company is contemplating the replacement of its old printing machine with a new model costing $ 6 0,000. The old machine, which originally cost $40,000, has 6 years of expected life remaining and a current book value of $ 25 ,000 versus a current market value of $ 14 ,000. Target's corporate tax rate is 40 percent. If Target sells the old machine at market value, what is the initial after-tax outlay for the new printing machine? Cash outflow must be a negative number! Round it a whole dollar and do not include the $ sign.
Answer #1

INITIAL AFTER TAX OUTLAY FOR NEW MACHINE = COST OF NEW MACHINE + AFTER TAX SALVAGE OLD MACHINE

THE BOOK VALUE OF OLD MACHINE = 25000

THE MARKET VALUE OF OLD MACHINE = 14000

SO THERE IS A LOSS OF 25000 - 14000 = 11000 ON SALE OF OLD MACHINE

TAX CREDIT ON LOSS ON SALE OF OLD MACHINE = 40% X 11000 = 4400

SO AFTER TAX SALVAGE VALUE = MARKET VALUE + TAX CREDIT ON LOSS ON SALE

AFTER TAX SALVAGE VALUE = 14000 + 4400 = 18400

INITIAL AFTER TAX OUTLAY FOR NEW MACHINE = -60000 + 18400 = -41600

Answer : - 41600 (Thumbs up please)